OpenStudy (anonymous):

c+(4-3c)-2=0 help me solve this plez

OpenStudy (nubeer):

first open brackets.. be careful with sign outside the bracket.. it would look like c+4-3c-2 = 0 right?

OpenStudy (anonymous):

ok yeah

OpenStudy (nubeer):

ok good.. now add the terms the c terms will only add or subtract with c. and the terms without c would be added or subtract speperately.

OpenStudy (nubeer):

c-3c +4 -2 = 0 just reaaranged it.. got it?

OpenStudy (nubeer):

tell me what will you get now?

OpenStudy (anonymous):

ok well I added 2 with 0 so I got c-3c+4=2

OpenStudy (anonymous):

so im guessing now I subtract 4 with 2?

OpenStudy (nubeer):

yes keep going.